# Robust Design and Evaluation of

### v5

@@ 2.2 Identification
In order to partially identify the conditional likelihood, expected welfare of a decision rule, and evaluation estimands, we consider a general class of restrictions on the extent to which the outcome can differ on average between unselected and selected units.
Define the confounding function $\delta(x):=\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$, which summarizes the average level of unobserved confounding conditional on the covariates. Our main assumption is that the confounding function satisfies
$\underline{\delta}(x;\eta)\leq\delta(x)\leq\overline{\delta}(x;\eta)$ (6)
@@
for all $x\in\mathcal{X}$ and some researcher-specified bounding functions $\underline{\delta}(\cdot;\eta)$, $\overline{\delta}(\cdot;\eta)$ with nuisance parameters $\eta$. Let $\Delta$ denote the set of all confounding functions satisfying Equation (6).
Equation (6) nests unconfoundedness as the special case with $\underline{\delta}(x;\eta)=\overline{\delta}(x;\eta)=0$ for all $x\in\mathcal{X}$ and the assumption-free, worst-case bounds with $\underline{\delta}(x;\eta)=-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$, $\overline{\delta}(x;\eta)=1-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ (e.g., Manski(89); Manski(94)-SampleSelectionBounds). Equation (6) interpolates between these extreme cases by restricting the extent to which the average outcome among unselected units can differ from the average outcome among selected units with the same covariates. As we show in the next section, alternative choices of bounding functions formalize common heuristics for dealing with missing data in applied research.
Under Equation (6), the evaluation estimands, conditional likelihood, and expected welfare of a decision rule are partially identified. Towards this, the conditional probability can be rewritten as $\mu^{*}(x)=\mu_{1}(x)+\pi_{0}(x)\delta(x)$. The identified set is the set of all values consistent with the bounds on the confounding function,
@@
The identified sets for the evaluation estimands, $\mathcal{H}(\mbox{perf}(s;\beta))$ and $\mathcal{H}(\mbox{perf}_{+}(s;\beta))$, and the expected welfare of a decision rule, $\mathcal{H}(U(d))$, are defined analogously. The identified sets are closed intervals. [^10] We focus on positive class performance $\mbox{perf}_{+}(s;\beta)$ since negative class performance $\mbox{perf}_{-}(s;\beta)$ is identical. In Online Appendix C.5, we derive the identified set for the difference in predictive performance across sensitive attributes under Equation (6), which can be used to evaluate a predictive algorithm’s fairness properties. As shorthand, let $\beta_{0,i}:=\beta_{0}(X_{i};s),\beta_{1,i}=\beta_{1}(X_{i};s)$ and $\underline{\delta}_{i}:=\underline{\delta}(X_{i};\eta)$, $\overline{\delta}_{i}:=\overline{\delta}_{i}$.
@@
$\mathcal{H}(\mbox{perf}(s;\beta))=\left[\underline{\mbox{perf}}(s;\beta),\overline{\mbox{perf}}(s;\beta)\right],$
$\mathcal{H}(\mbox{perf}_{+}(s;\beta))=\left[\underline{\mbox{perf}}_{+}(s;\beta),\overline{\mbox{perf}}_{+}(s;\beta)\right],$
@@
$\overline{\mbox{perf}}_{+}(s;\beta)=\sup_{\delta(\cdot)\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})],$
$\underline{\mbox{perf}}_{+}(s;\beta)=\inf_{\delta(\cdot)\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})].$
@@
For $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ and $u_{d,y^{*},i}=u_{d,y^{*}}(X_{i})$, $\mathcal{H}(U(d))=[\underline{U}(d),\overline{U}(d)]$, where
@@ 2.3 Choice of bounding functions
The bounding functions $\underline{\delta}(\cdot;\eta)$, $\overline{\delta}(\cdot;\eta)$ in Equation (6) must be specified by the researcher, and the choice will depend on the empirical application at hand. Through alternative choices of these bounding functions, our framework formalizes popular strategies for addressing unobserved confounding in empirical applications. We next highlight several examples.
2.3.1 Nonparametric outcome bounds
In some empirical settings, it may be reasonable to assume the unobserved probability of $Y_{i}^{*}=1$ among unselected units cannot be too different than the observed probability of $Y_{i}^{*}=1$ among selected units. This can be formalized as
$\underline{\delta}(x;\eta):=(\underline{\Gamma}-1)\mu_{1}(x)\mbox{ and }\overline{\delta}(x;\eta):=(\overline{\Gamma}-1)\mu_{1}(x)$ (7)
for all $x\in\mathcal{X}$. This choice implies $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)$ is bounded above by $\overline{\Gamma}\mu_{1}(x)$ and below by $\underline{\Gamma}\mu_{1}(x)$. Unconfoundedness is now nested as the special case with $\overline{\Gamma}=\underline{\Gamma}=1$. We refer to this choice as nonparametric outcome bounds.
Nonparametric outcome bounds formalize popular heuristics for imputing missing data based on observed outcomes among selected units. For example, KLLLM(18) evaluates an estimated pretrial risk assessment tool under varying assumptions on how much the average failure to appear rate among detained defendants may vary relative to released defendants (see their Table 5). In consumer finance, researchers apply coarse adjustments to the observed default rate among accepted applicants to impute the missing default rate among rejected applicants (e.g., HandHenley(93); ZengZhao(14)). By specifying the bounding functions to be proportional to the observed conditional probability, nonparametric outcome bounds flexibly impute the missing data in a manner that richly depends on the covariates.
Remark 1 (Connection to other sensitivity analysis models in causal inference).
In causal inference, there exists several alternative relaxations of unconfoundedness, such as the marginal sensitivity model (e.g., Tan(06); DornGuo(21)), Rosenbaum’s sensitivity model (e.g., Rosenbaum(02); Yadlowsky(21)evaluating), and partial $c$-dependence (e.g., MastenPoirier(18); MastenPoirier(20)). Unlike our framework, these sensitivity analysis models specify bounds on how unobserved confounders may have affected historical selection decisions, which may be difficult to reason about when past decisions were the choices of human decision makers. Nonetheless, as we show in Online Appendix D, each of these sensitivity analysis models imply nonparametric outcome bounds with particular choices of $\underline{\Gamma},\overline{\Gamma}$ in our setting. Our analysis of nonparametric outcome bounds therefore immediately applies to these sensitivity analysis models.
@@ 2.3.2 Proxy outcome bounds
In other empirical settings, we may observe an additional proxy outcome $\tilde{Y}_{i}\in\{0,1\}$ that is statistically related to $Y_{i}^{*}$ and always observed. In consumer finance, we observe whether the applicant defaulted on other approved credit products like a credit card, and this may serve as a proxy outcome for the default outcome of interest on a mortgage (BlattnerNelson(21)). In health care, we may use a patient’s longer term health outcomes as a proxy for whether they suffered from the medical condition (ChanGentzkowYu(21); MullainathanObermeyer(21)), or their total medical spending as a proxy for their underlying health status (ObermeyerEtAl(19)).
We may use the observed proxy outcome $\tilde{Y}_{i}$ to impute missing outcomes by assuming that the proxy outcome is as informative about true outcome among unselected units as it is among selected units. This can be formalized as a particular choice of bounding functions in our framework. Suppose the proxy outcome satisfies, for all $x\in\mathcal{X}$,
$\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=0,X_{i}=x)=\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x).$ (8)
Under this assumption, the confounding function satisfies Equation (6) with bounding functions
$\underline{\delta}(x;\eta)=1-\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)-\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)-\mu_{1}(x),$
$\overline{\delta}(x;\eta)=1-\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)+\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)-\mu_{1}(x),$
provided $\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)\leq\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)$ and $\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X=x)+\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)\leq 1$ for all $x\in\mathcal{X}$. These restrictions on the proxy outcome simplify the bounding functions and are satisfied on average by, for example, the proxy outcomes used in BlattnerNelson(21) and MullainathanObermeyer(21). In Online Appendix C.3, we provide a more general characterization of the bounding functions based on proxy outcomes without these additional restrictions. To our knowledge, there is no analogue of such proxy outcome assumptions in the existing causal inference literature.
@@ 2.3.3 Instrumental variable bounds
In some empirical settings, we may observe an instrumental variable $Z_{i}\in\mathcal{Z}$ with finite support that generates variation in the historical decision and satisfies $Y_{i}^{*}\perp\!\!\!\perp Z_{i}\mid X_{i}$. A popular example arises through the random assignment of historical decision makers to cases, such as judges to defendants in the pretrial release system (e.g., KLLLM(18); ArnoldDobbieHull(20)-discrimination_bail; Rambachan(21); AngelovaEtAl(22)) or hiring managers to applicants (LiRaymondBergman(20)). Other empirical examples include variation in the shift or day-of-week that a patient arrives to the hospital (e.g., MullainathanObermeyer(21)) or idiosyncratic timing variation in banking regulations across U.S. states (e.g., BlattnerNelson(21)).
Such an instrument implies particular bounding functions in our framework. Defining $\pi_{d}(x,z):=\mathbb{P}(D_{i}=d\mid X_{i}=x,Z_{i}=z)$, the confounding function satisfies, for any $z\in\mathcal{Z}$,
$\underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\overline{\delta}_{z}(x;\eta)$
for all $x\in\mathcal{X}$ and $\underline{\delta}_{z}(x)=\left(\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]-\mu_{1}(x)\right)/\pi_{0}(x)$ and $\overline{\delta}_{z}(x):=(\pi_{0}(x,z)+\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]-\mu_{1}(x))/\pi_{0}(x)$. This is an immediate consequence of results in Manski(94)-SampleSelectionBounds. For any fixed value $z\in\mathcal{Z}$, we can apply our identification and estimation results based on Equation (6).
We may further aggregate across all values of the instrument to obtain sharper bounds of the form
These bounds are pointwise intersection bounds (Chernozhukov-IntersectionBounds). Due to the maximum and minimum operators, however, these bounding functions are non-smooth without further assumptions, posing a technical challenge for our estimation procedures that rely on a debiasing step based on efficient influence functions. We instead suggest bounding functions based on smooth approximations to these intersection bounds, extending LevisEtAl(23)-CovariateAssistedIV’s analysis of the covariate-assisted Balke-Pearl bounds on the average treatment effect to the large class of predictive estimands and discrete instruments. [^11] LevisEtAl(23)-CovariateAssistedIV and semenova2023adaptive consider an alternative strategy for dealing with pointwise intersection bounds that invokes a “margin condition,” which states the maximal or minimal values of the intersection bounds are sufficiently well-separated. We leave the analysis of margin conditions in our setting for future work. We discuss both instrumental variable bounds based on a fixed value $z\in\mathcal{Z}$ and smooth approximations to these intersection bounds in Online Appendix C.4.

### v6

@@ 2.2 Identification
Selective labels create an identification problem: since $Y_{i}^{*}$ is only observed when $D_{i}=1$, the conditional likelihood $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i})$ is not point identified from the data alone. Consequently, neither are the conditional likelihood, expected payoffs for any decision rule, nor predictive performance measures defined in Section 2.1.
A natural approach would be to assume unconfounded selection — that is, $Y_{i}^{*}\perp\!\!\!\perp D_{i}\mid X_{i}$ (e.g., Coston(20)-Evaluation; mishler2021fade). Under unconfoundedness, the conditional likelihood among unselected units equals that among selected units with the same covariates: $P(Y_{i}^{*}=1\mid D_{i}=0,X_{i})=P(Y_{i}^{*}=1\mid D_{i}=1,X_{i})$. However, this assumption is often implausible in practice. Selected and unselected units may differ systematically in ways not captured by observed covariates $X_{i}$. In consumer lending, applicants’ decisions to accept loan terms may depend on competing offers from other lenders or their private assessment of their ability to repay — information unavailable to the researcher. In pretrial release, judges may rely on courtroom interactions or details from case files that are not recorded in administrative data. Consequently, unobserved confounding — unobservables correlated with both the outcome $Y_{i}^{*}$ and the decision $D_{i}$ — is a central challenge.
Rather than imposing unconfoundedness, we partially identify the conditional likelihood, expected payoffs, and predictive performance measures by placing restrictions on how much outcomes can differ between unselected and selected units with the same covariates. These restrictions formalize researcher beliefs about the severity of unobserved confounding in a transparent way and nest popular empirical strategies as special cases. The key to our approach is bounding the confounding function $\delta(x):=\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$, which summarizes the difference in outcomes between unselected and selected units conditional on covariates. When $\delta(x)=0$ for all $x\in\mathcal{X}$, the selection decision is unconfounded. When $\delta(x)\neq 0$, unobserved confounding is present: units with the same observed covariates have systematically different outcomes depending on whether they were selected.
Assumption 2.1.
@@
For researcher-specified bounding functions $\underline{\delta}(\cdot;\eta),\overline{\delta}(\cdot;\eta)$ with (possibly infinite-dimensional) nuisance parameters $\eta$, the confounding function satisfies
$\underline{\delta}(x;\eta)\leq\delta(x)\leq\overline{\delta}(x;\eta)\mbox{ for all }x\in\mathcal{X}.$ (4)
Let $\Delta$ denote the set of all confounding functions satisfying Equation (4).
Assumption 2.1 restricts how much outcomes can differ between unselected and selected units with the same observed covariates. In consumer lending, this bounds the difference in default rates between unfunded and funded applicants conditional on application characteristics. In pretrial release, this bounds the difference in misconduct rates between detained and released defendants conditional on charge and criminal history.
Assumption 2.1 allows researchers to encode beliefs about unobserved confounding without committing to a specific selection model. It nests two important special cases: unconfoundedness with $\underline{\delta}(x;\eta)=\overline{\delta}(x;\eta)=0$ for all $x\in\mathcal{X}$, and the assumption-free, worst-case bounds with $\underline{\delta}(x;\eta)=-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$, $\overline{\delta}(x;\eta)=1-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ (e.g., Manski(89); Manski(94)-SampleSelectionBounds). Thus Assumption 2.1 interpolates between strong identifying assumptions and possibly uninformative bounds.
More broadly, Assumption 2.1 follows the structure of widely-used partial identification frameworks in econometrics, most notably Manski(03)’s ”approximate mean independence” assumption. It extends this by conditioning on covariates and allowing bounds to depend on identified nuisance parameters. In this sense, Assumption 2.1 formalizes restrictions in terms that align with domain knowledge: researchers in these settings naturally reason about outcome differences rather than, say, how unobservables affect selection propensities. As shown in Section 2.3, the bounding functions can be chosen through direct calibration based on domain expertise or derived indirectly from additional data features such as proxy outcomes or instrumental variables. Each strategy is instantiated by specifying both the functional form of the bounding functions and the collection of nuisance parameters $\eta$ (possibly infinite-dimensional) to be estimated.
Under Assumption 2.1, the conditional likelihood and evaluation estimands are partially identified. We assume throughout that the chosen bounding functions yield valid conditional probabilities. Specifically, for all $\delta\in\Delta$ and $x\in\mathcal{X}$, $\mu_{1}(x)+\pi_{0}\delta(x;\eta)\in[0,1]$ and $\delta(x;\eta)\in[-1,1]$. This is a mild restriction in practice; as we discuss below, for common bounding functions, these constraints can be enforced through the researcher’s specification of sensitivity parameters or are satisfied automatically, as we will discuss next. If these constraints are violated and the researcher proceeds with our identification framework, the resulting bounds will be conservative (non-sharp) but may still be informative.
Since $\mu^{*}(x)=\mu_{1}(x)+\pi_{0}(x)\delta(x)$, the identified set is the set of all values consistent with the bounds on $\delta(x)$,
@@
The identified sets for the performance measures, $\mathcal{H}(\mbox{perf}(s;\beta))$ and $\mathcal{H}(\mbox{perf}_{+}(s;\beta))$, and for the expected payoff of a decision rule $\mathcal{H}(U(d))$ are defined analogously. As shorthand, let $\beta_{0,i}:=\beta_{0}(X_{i};s),\beta_{1,i}=\beta_{1}(X_{i};s)$ and $\underline{\delta}_{i}:=\underline{\delta}(X_{i};\eta)$, $\overline{\delta}_{i}:=\overline{\delta}_{i}(X_{i};\eta)$.
@@
[…]
@@
$\overline{\mbox{perf}}_{+}(s;\beta)=\sup_{\delta\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})],$
$\underline{\mbox{perf}}_{+}(s;\beta)=\inf_{\delta\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})].$
@@
For $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ and $u_{d,y^{*},i}=u_{d,y^{*}}(X_{i})$, $\mathcal{H}(U(d))=[\underline{U}(d),\overline{U}(d)]$ for $\underline{U}(d),\overline{U}(d)$ defined in the proof.
The remainder of the paper applies this partial identification framework to the two objectives for predictive algorithms commonly seen in applied work. Section 3 addresses algorithm design: we estimate bounds on the conditional likelihood, thereby constructing predictive algorithms robust to unobserved confounding. Section 4 addresses algorithm evaluation: we estimate bounds on performance measures for a given predictive algorithm. Before proceeding to these estimation results, we next discuss how researchers might choose the bounding functions in Assumption 2.1.
Remark 1 (Connection to Algorithmic Fairness).
Differences in predictive performance across sensitive attributes (e.g., race, gender) formalize violations of popular fairness criteria studied in the algorithmic fairness literature (see MitchellEtAl(19); barocas-hardt-narayanan, for reviews). Specifically, differences in positive and negative class performance across groups characterize violations of “equality of opportunity” (hardt2016equality), “equalized odds” (chouldechova2017fair; KleinbergMullainathanRaghavan(17)), and “group-level discrimination” (ArnoldDobbieHull(20)-discrimination_alg; ArnoldDobbieHull(20)-discrimination_bail). Differences in overall performance across groups characterize violations of “bounded group loss” (AgarwalEtAl(19)-FairRegression). In Online Appendix C.1, we partially identify these performance disparities under Assumption 2.1, enabling robust fairness audits.
@@ 2.3 Choice of Bounding Functions
How should the bounding functions $\underline{\delta}(\cdot;\eta)$ and $\overline{\delta}(\cdot;\eta)$ be chosen? The choice encodes assumptions about how much unselected units might differ from selected units with the same observed characteristics. In consumer lending: How much more likely are unfunded applicants to default than funded applicants with the same credit profile? In pretrial release: How much more likely are detained defendants to commit misconduct than released defendants with the same criminal history? We next show how alternative bounding functions formalize three common strategies for addressing selective labels in applied research.
2.3.1 Observed Outcome Bounds
A natural starting point is to assume that the unobserved outcome rate among unselected units is similar to the observed outcome rate among selected units with the same characteristics. For instance, in pretrial release, detained defendants with a given criminal history might be 1.5 times as likely to commit misconduct as released defendants with the same history. In consumer lending, unfunded applicants might be twice as likely to default as funded applicants with the same credit profile.
Observed outcome bounds formalize this intuition by specifying, for researcher-chosen constants $\underline{\Gamma},\overline{\Gamma}>0$ and all $x\in\mathcal{X}$,
[…]
Under this choice, the unobserved conditional probability among unselected units satisfies $\underline{\Gamma}\mu_{1}(x)\leq P(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)\leq\overline{\Gamma}\mu_{1}(x)$ — that is, it is bounded by constant multiples of the observed conditional probability $\mu_{1}(x)$. Setting $\underline{\Gamma}=\overline{\Gamma}=1$ recovers unconfoundedness. For observed outcome bounds, the probability constraints are satisfied for all choices of $\underline{\Gamma},\overline{\Gamma}$ in a neighborhood of one under strict overlap on the decision and observed outcome probabilities being bounded away from zero and one.
This approach formalizes strategies commonly used in applied work. KLLLM(18) evaluate a constructed pretrial risk tool under varying assumptions on how much misconduct rates among detained defendants differ from released defendants (their Table 5). In consumer finance, researchers impute default rates for rejected applicants by applying multiplicative adjustments to observed default rates among accepted applicants (e.g., HandHenley(93); ZengZhao(14)). Observed outcome bounds flexibly encode such adjustments while allowing the imputed values to vary richly with covariates.
Remark 2 (Connection to Sensitivity Analysis Models in Causal Inference).
Several influential sensitivity analysis frameworks — including the marginal sensitivity model (Tan(06); ZhaoSmallBhattacharya(19); DornGuo(21)), Rosenbaum’s $\Gamma$-sensitivity model (Rosenbaum(02); Yadlowsky(21)evaluating), and partial $c$-dependence (MastenPoirier(18); MastenPoirier(20)) — place bounds on how unobservables affect selection propensities. While conceptually distinct from bounding outcome differences, these models imply observed outcome bounds with specific choices of $\underline{\Gamma},\overline{\Gamma}$ (see Online Appendix E). The connection between outcome-based bounds and selection-based sensitivity models arises through Bayes’ rule; though the observed outcome bounds we derive may not exhaust all implications of the original sensitivity model.
@@ 2.3.2 Proxy Outcome Bounds
Sometimes researchers observe a proxy outcome $\widetilde{Y}_{i}\in\{0,1\}$ for both selected and unselected units that is statistically related to the true outcome $Y_{i}^{*}$. This proxy can help discipline assumptions about unobserved confounding. For example, in consumer lending, whether an applicant defaulted on other credit products (e.g., credit cards) serves as a proxy for whether they would default on the loan of interest (BlattnerNelson(21)). In medical testing, long-term health outcomes or total medical spending serves as a proxy for whether a patient suffered an acute condition (ObermeyerEtAl(19); ChanGentzkowYu(21); MullainathanObermeyer(21)).
The key assumption is that the proxy outcome is equally informative about the true outcome for both selected and unselected units. Formally, suppose the proxy outcome satisfies, for all $x\in\mathcal{X}$,
$P(Y^{*}_{i}=\tilde{Y}_{i}\mid D_{i}=0,X_{i}=x)=P(Y^{*}_{i}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x).$ (6)
Conditional on the covariates, the probability that the true and proxy outcome agree is the same regardless of selection status. This assumption, combined with the observed distribution of $\widetilde{Y}_{i}$ among unselected units, pins down specific bounding functions on $\delta(x)$.
Under mild restrictions satisfied by the proxy outcomes used in BlattnerNelson(21) and MullainathanObermeyer(21), the bounding functions simplify to
$\underline{\delta}(x;\eta)=1-\gamma_{1}(x)-\widetilde{\mu}_{0}(x)-\mu_{1}(x),$
$\overline{\delta}(x;\eta)=1-\gamma_{1}(x)+\widetilde{\mu}_{0}(x)-\mu_{1}(x),$
where $\gamma_{1}(x):=\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)$ is the probability the true and proxy outcome agree among selected units and $\widetilde{\mu}_{0}(x):=\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)$ is the observed proxy outcome rate among unselected units. These bounding functions satisfy the probability constraints under the maintained proxy assumptions under strict overlap on the decision and observed outcome probabilities being bounded away from zero and one. Online Appendix D.1 provides the general characterization without these simplifying restrictions. To our knowledge, proxy outcome assumptions of this form have not been explored in the causal inference sensitivity analysis literature.
@@ 2.3.3 Instrumental Variable Bounds
Sometimes researchers observe an instrumental variable $Z_{i}\in\mathcal{Z}$ with finite support that generates quasi-random variation in selection decisions but does not directly affect outcomes (i.e., $Y_{i}^{*}$ independent of $Z_{i}$ given $X_{i}$). For example, in pretrial release, judges randomly assigned to cases vary in their propensity to release defendants (KLLLM(18); ArnoldDobbieHull(20)-discrimination_bail; Rambachan(21); AngelovaEtAl(22)). The day-of-week affects whether certain tests are ordered (e.g., MullainathanObermeyer(21)), and state-level variation in banking regulations affects loan approval rates (e.g., BlattnerNelson(21)).
An instrument $Z_{i}$ implies specific bounding functions through a classic result of Manski(94)-SampleSelectionBounds. For any fixed instrument value $z\in\mathcal{Z}$, the confounding function satisfies, for all $x\in\mathcal{X}$,
where $\underline{\delta}_{z}(x)=\left(\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]-\mu_{1}(x)\right)/\pi_{0}(x)$ and $\overline{\delta}_{z}(x):=(\pi_{0}(x,z)+\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]-\mu_{1}(x))/\pi_{0}(x)$. Our identification and estimation results apply to these bounds for any fixed $z\in\mathcal{Z}$.
Researchers can obtain sharper bounds by aggregating across all instrument values: $\max_{z\in\mathcal{Z}}\ \underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\min_{z\in\mathcal{Z}}\ \overline{\delta}_{z}(x;\eta)$. These intersection bounds (Chernozhukov-IntersectionBounds) are tighter but involve maximum and minimum operators, creating non-smoothness that complicates estimation. Following LevisEtAl(23)-CovariateAssistedIV, we address this challenge using smooth approximations to the intersection bounds. Online Appendix D.2 discusses both approaches: bounds based on a fixed instrument value and smooth approximations to intersection bounds. [^6] LevisEtAl(23)-CovariateAssistedIV and semenova2023adaptive consider an alternative strategy for dealing with pointwise intersection bounds that invokes a “margin condition” on the separation of the maximal value of the intersection bounds. Under such a margin condition, existing results would apply to bounds on overall performance. Extending this approach to estimators for the conditional likelihood bounds or positive class performance bounds is not immediate. We leave the full analysis of margin conditions for future work.